# Online Convex Optimization with Heavy Tails: Old Algorithms, New Regrets, and Applications

### v1

@@ Abstract
In Online Convex Optimization (OCO), when the stochastic gradient has a finite variance, many algorithms provably work and guarantee a sublinear regret. However, limited results are known if the gradient estimate has a heavy tail, i.e., the stochastic gradient only admits a finite $\mathsf{p}$-th central moment for some $\mathsf{p}\in\left(1,2\right]$. Motivated by it, this work examines different old algorithms for OCO (e.g., Online Gradient Descent) in the more challenging heavy-tailed setting. Under the standard bounded domain assumption, we establish new regrets for these classical methods without any algorithmic modification. Remarkably, these regret bounds are fully optimal in all parameters (can be achieved even without knowing $\mathsf{p}$), suggesting that OCO with heavy tails can be solved effectively without any extra operation (e.g., gradient clipping). Our new results have several applications. A particularly interesting one is the first provable convergence result for nonsmooth nonconvex optimization under heavy-tailed noise without gradient clipping. Furthermore, we explore broader settings (e.g., smooth OCO) and extend our ideas to optimistic algorithms to handle different cases simultaneously.
@@ 1.1 Contributions
  • We extend the analysis of $\mathsf{OGD}$ to Online Strongly Convex Optimization with heavy tails and establish the first provable result $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\log T}{\mu}+\frac{\sigma^{\mathsf{p}}G^{2-\mathsf{p}}}{\mu}T^{2-\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}$, where $\mu>0$ is the modulus of strong convexity and $T^{0}$ should be read as $\log T$.
@@
Furthermore, we explore broader settings. For example, when each $\ell_{t}$ is $H$-smooth, we present new regrets that extend the classical $L^{\star}$ bounds [29] to the heavy-tailed noise case. As an important implication, we show that $\mathsf{SGD}$ converges at a rate of $HD^{2}/T+\sigma D/T^{1-1/\mathsf{p}}$ for smooth convex optimization with heavy tails. Finally, we extend our ideas to optimistic algorithms to address various cases simultaneously and employ optimistic algorithms to give the first provable result for Hölder smooth nonconvex optimization under heavy tails, where the problem-dependent parameters can be either known or unknown.
@@ 1.2 Discussion on [51]
Moreover, we would like to mention two drawbacks of [51]. First, though the $T^{1/\mathsf{p}}$ regret seems tight as it matches the lower bound [26, 32, 45], this may not be the best, since an optimal bound should recover the standard $\sqrt{T}$ regret in the deterministic case (i.e., $\sigma=0$), as one can imagine. This suggests that their bound is not entirely optimal. Second, we remark that they require knowing both problem-dependent parameters $G$, $\sigma$, $\mathsf{p}$ and time horizon $T$ in the algorithm, which may be hard to satisfy in the online setting. In comparison, our regret bound $GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}$ is fully optimal in all parameters. Importantly, $\mathsf{AdaGrad}$ can achieve it while oblivious to the problem information.
@@ 3.1 New Regret for Online Gradient Descent
A less well known analysis [^1] To clarify, the phrase “less well known” is compared to the first one. This analysis itself is also well known.. As discussed, the failure of the above proof under heavy-tailed noise is due to (1). Therefore, if a tighter inequality than (1) exists, then it might be possible to show that $\mathsf{OGD}$ still works for $\mathsf{p}<2$. However, does it exist?
@@ 3.2 New Regret for Dual Averaging
It is known that $\mathsf{DA}$ is a special realization of the more general Follow-the-Regularized-Leader $(\mathsf{FTRL})$ framework [24]. To keep the work concise, we only focus on $\mathsf{DA}$. The key idea to prove Theorem 2 can directly extend to show new regret for $\mathsf{FTRL}$ under heavy-tailed noise.
@@ 4.1 Nonsmooth Convex Optimization
Convergence of the average iterate. First, we focus on convergence in average. By the classical online-to-batch conversion [4], the following corollary immediately holds.
@@
We are only aware of two works [22, 45] based on Stochastic Mirror Descent ($\mathsf{SMD}$) [26] that gave convergence results without clipping. However, they share a common drawback, i.e., their bounds are both in the form of $(G+\sigma)D/T^{1-1/\mathsf{p}}$, which cannot recover the optimal rate $GD/\sqrt{T}$ when $\sigma=0$.
@@
Convergence of the last iterate. Next, we consider the more challenging last-iterate convergence, which has a long history in stochastic optimization and fruitful results in the case of $\mathsf{p}=2$ (see, e.g., [30, 39, 53]). However, less is known about heavy-tailed problems. So far, only two works [22, 31] have established the last-iterate convergence. The former is based on $\mathsf{SMD}$, and the latter employs gradient clipping in $\mathsf{SGD}$. Unfortunately, their rates are both in the suboptimal order $(G+\sigma)D/T^{1-1/\mathsf{p}}$.
@@
One may notice that $\boldsymbol{y}_{t}$ itself is not the decision made by the online learner and naturally may ask whether $\boldsymbol{x}_{t}$ ensures the last-iterate convergence if we simply pick $\ell_{t}=F$. The answer turns out to be positive at least for $\mathsf{OGD}$ (which is equivalent to $\mathsf{SGD}$ now). However, to prove this result, we rely on a technique specialized to stochastic optimization recently developed by [22, 48]. To not diverge from the topic of OCO, we defer the last-iterate convergence of $\mathsf{OGD}$ to Appendix D, in which Theorem 10 gives a general result for any stepsize $\eta_{t}$ and Corollary 5 shows a last-iterate rate similar to Corollary 2 (up to an extra logarithmic factor) under the same stepsize $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ as in Theorem 1.
@@ 4.2 Nonsmooth Nonconvex Optimization
However, on the theoretical side, their result cannot recover the optimal bound $G^{2}\delta^{-1}\epsilon^{-3}$ [7] in the deterministic case. On the practical side, their method also employs the gradient clipping step, which introduces a new clipping parameter to tune. In fact, as stated in their Section 5, they observed in experiments that their algorithm without the clipping operation (exactly the algorithm we study next) still works under heavy tails. In addition, in their Section 6, they also explicitly ask whether the requirement to know G and A can be removed.
@@ 5.2 Optimistic Algorithms for Broader Cases
For example, in the special case of heavy-tailed Hölder smooth optimization (i.e., $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu}$ for some $H>0$ and $\nu\in\left(0,1\right]$), Corollary 8 provides a rate $\frac{HD^{1+\nu}}{T^{\frac{1+\nu}{2}}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$ for convex problems, Corollaries 11 and 12 establish provable results for nonconvex problems, where the problem-dependent parameters may be known or unknown. All of these results are new to the best of our knowledge.
6 Conclusion and Limitation
@@ A.2 Extension to Online Strongly Convex Optimization
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}G^{2-\mathsf{p}}}{\mu}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases},\forall\boldsymbol{x}\in\mathcal{X}.$
@@
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]$ $\lesssim\sum_{t=1}^{T}\frac{G^{2}}{\mu t}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}t^{\mathsf{p}-1}}$
$\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases}.$
Lastly, it is known that if $\ell_{t}$ is $G$-Lipschitz and $\mu$-strongly convex on a domain $\mathcal{X}$ with a diameter $D$, then it satisfies $D\lesssim\frac{G}{\mu}$ (e.g., see Lemma 2 of [34]). Therefore, when $\mathsf{p}\in\left(1,2\right)$,
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}G^{2-\mathsf{p}}}{\mu}T^{2-\mathsf{p}}.$
@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
We prove the following last-iterate convergence result for $\mathsf{SGD}$ (i.e., $\mathsf{OGD}$ for stochastic optimization) under heavy-tailed noise. The proof of Theorem 10 is inspired by [22, 48].
@@
Under Assumption 1 for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, for any stepsize $\eta_{t}>0$ in $\mathsf{OGD}$ (Algorithm 1), we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{D^{2}}{\sum_{t=1}^{T}\eta_{t}}+G^{2}\sum_{t=1}^{T}\frac{\eta_{t}^{2}}{\sum_{s=(t+1)\land T}^{T}\eta_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T}\frac{\eta_{t}^{\mathsf{p}}}{\sum_{s=(t+1)\land T}^{T}\eta_{s}},\forall\boldsymbol{x}\in\mathcal{X}.$
@@
$w_{t}\triangleq\frac{\eta_{T}}{\sum_{s=t+1}^{T}\eta_{s}},\forall t\in\left\{0\right\}\cup\left[T-1\right]$ and $w_{T}\triangleq w_{T-1}=1.$ (27)
@@
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}}{2\eta_{t}}+\eta_{t}G^{2}+\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (29)

### v2

@@ Abstract
In Online Convex Optimization (OCO), when the stochastic gradient has a finite variance, many algorithms provably work and guarantee a sublinear regret. However, limited results are known if the gradient estimate has a heavy tail, i.e., the stochastic gradient only admits a finite $\mathsf{p}$-th central moment for some $\mathsf{p}\in\left(1,2\right]$. Motivated by it, this work examines different old algorithms for OCO (e.g., Online Gradient Descent) in the more challenging heavy-tailed setting. Under the standard bounded domain assumption, we establish new regrets for these classical methods without any algorithmic modification. Remarkably, these regret bounds are fully optimal in all parameters (can be achieved even without knowing $\mathsf{p}$), suggesting that OCO with heavy tails can be solved effectively without any extra operation (e.g., gradient clipping). Our new results have several applications. A particularly interesting one is the first provable and optimal convergence result for nonsmooth nonconvex optimization under heavy-tailed noise without gradient clipping. Furthermore, we explore broader settings (e.g., smooth OCO) and extend our ideas to optimistic algorithms to handle different cases simultaneously.
@@ 1.1 Contributions
  • We extend the analysis of $\mathsf{OGD}$ to Online Strongly Convex Optimization with heavy tails and establish the first provable result $\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\log T}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}T^{2-\mathsf{p}},\forall\boldsymbol{x}\in\mathcal{X}$, where $\mu>0$ is the modulus of strong convexity and $T^{0}$ should be read as $\log T$.
@@
Furthermore, we explore broader settings. For example, when each $\ell_{t}$ is $H$-smooth, we present new regrets that extend the classical $L^{\star}$ bounds [32, 43] to the heavy-tailed noise case. As an important implication, we show that $\mathsf{SGD}$ converges at a rate of $HD^{2}/T+\sigma D/T^{1-1/\mathsf{p}}$ for smooth convex optimization with heavy tails (even for the last iterate). Finally, we extend our ideas to optimistic algorithms to address various cases simultaneously and employ optimistic algorithms to give the first provable result for Hölder smooth nonconvex optimization under heavy tails, where the problem-dependent parameters can be either known or unknown.
@@ 1.2 Discussion on [53]
Moreover, we would like to mention two limitations of [53]. First, though the $T^{1/\mathsf{p}}$ regret seems tight as it matches the lower bound [29, 35, 47], this may not be the best, since an optimal bound should recover the standard $\sqrt{T}$ regret in the deterministic case (i.e., $\sigma=0$), as one can imagine. This suggests that their bound is not entirely optimal. Second, we remark that they require knowing both problem-dependent parameters $G$, $\sigma$, $\mathsf{p}$ and time horizon $T$ in the algorithm, which may be hard to satisfy in the online setting. In comparison, our regret bound $GD\sqrt{T}+\sigma DT^{1/\mathsf{p}}$ is fully optimal in all parameters. Importantly, $\mathsf{AdaGrad}$ can achieve it while oblivious to the problem information.
@@ 2 Preliminary
Remark 3.
The idea presented in this work can be extended to the composite setting (i.e., $\ell_{t}+\psi_{t}$, where $\psi_{t}$ is convex and assumed to be known at time t1t-1) via the proximal update, provably.
@@ 3.1 New Regret for Online Gradient Descent
A less well known analysis [^1] To clarify, the phrase “less well known” is compared to the first one. This analysis itself is also well known [22, 32].. As discussed, the failure of the above proof under heavy-tailed noise is due to (1). Therefore, if a tighter inequality than (1) exists, then it might be possible to show that $\mathsf{OGD}$ still works for $\mathsf{p}<2$. However, does it exist?
@@ 3.2 New Regret for Dual Averaging
It is known that $\mathsf{DA}$ is a special realization of the more general Follow-the-Regularized-Leader $(\mathsf{FTRL})$ framework [27]. To keep the work concise, we focus only on $\mathsf{DA}$. The key idea for proving Theorem 2 can be directly extended to show new regret for $\mathsf{FTRL}$ under heavy-tailed noise.
@@ 4.1 Nonsmooth Convex Optimization
Convergence of the average iterate. First, we focus on the average-iterate convergence. By the classical online-to-batch conversion [4], the following corollary holds.
@@
We are only aware of two works [25, 47] based on Stochastic Mirror Descent ($\mathsf{SMD}$) [29] that gave convergence results without clipping. However, they share a common shortcoming, i.e., their bounds are both in the form of $(G+\sigma)D/T^{1-1/\mathsf{p}}$, which cannot recover the optimal rate $GD/\sqrt{T}$ when $\sigma=0$.
@@
Convergence of the last iterate. Next, we consider the more challenging last-iterate convergence, which has a long history in stochastic optimization and fruitful results in the case of $\mathsf{p}=2$ (see, e.g., [11, 16, 33, 41, 55]). However, less is known about heavy-tailed problems. So far, only two works [25, 34] have established the last-iterate convergence. The former is based on $\mathsf{SMD}$, and the latter employs gradient clipping in $\mathsf{SGD}$. Unfortunately, their rates are both in the suboptimal order $(G+\sigma)D/T^{1-1/\mathsf{p}}$.
@@
One may notice that $\boldsymbol{y}_{t}$ itself is not the decision made by the online learner and naturally may ask whether $\boldsymbol{x}_{t}$ ensures the last-iterate convergence if we simply pick $\ell_{t}=F$. The answer turns out to be positive at least for $\mathsf{OGD}$ (which is equivalent to $\mathsf{SGD}$ now). However, to prove this result, we rely on a technique specialized to stochastic optimization recently developed by [25, 50]. To not diverge from the topic of OCO, we defer the last-iterate convergence of $\mathsf{OGD}$ (i.e., $\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})\right]$) to Appendix D, in which Theorem 10 gives a general result for any stepsize $\eta_{t}$ (even for a broader function class) and Corollary 5 shows a last-iterate rate similar to Corollary 2 (up to an extra logarithmic factor) under the same stepsize $\eta_{t}=\frac{D}{G\sqrt{t}}\land\frac{D}{\sigma t^{1/\mathsf{p}}}$ as in Theorem 1. If $T$ is assumed to be known, Corollary 5 also shows how to set the stepsize in $\mathsf{OGD}$ to converge in the optimal rate $\frac{GD}{\sqrt{T}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$. If strong convexity is additionally assumed, Corollary 6 gives a last-iterate rate in the order of $\frac{\log T}{T^{\mathsf{p}-1}}$.
@@ 4.2 Nonsmooth Nonconvex Optimization
However, on the theoretical side, their result cannot recover the optimal bound $G^{2}\delta^{-1}\epsilon^{-3}$ [7] in the deterministic case. On the practical side, their method also employs the gradient clipping step, which introduces a new clipping parameter to tune. In fact, as stated in their Section 5, they observed in experiments that their algorithm without the clipping operation (exactly the algorithm we study next) still works under heavy tails. In addition, in their Section 6, they also explicitly ask whether the requirement to know $G$ and $\mathsf{A}$ can be removed.
@@ 5.2 Optimistic Algorithms for Broader Cases
For example, in the special case of heavy-tailed Hölder smooth optimization (i.e., $\left\|\nabla F(\boldsymbol{x})-\nabla F(\boldsymbol{y})\right\|\leq H\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{\nu}$ for some $H>0$ and $\nu\in\left(0,1\right]$), Corollary 9 provides a rate $\frac{HD^{1+\nu}}{T^{\frac{1+\nu}{2}}}+\frac{\sigma D}{T^{1-\frac{1}{\mathsf{p}}}}$ for convex problems, Corollaries 12 and 13 establish provable results for nonconvex problems, where the problem-dependent parameters may be known or unknown. All of these results are new to the best of our knowledge.
6 Conclusion and Future Work
@@ A.2 Extension to Online Strongly Convex Optimization
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases},\forall\boldsymbol{x}\in\mathcal{X}.$
@@
$\mathbb{\mathbb{E}}\left[\mathsf{R}_{T}^{\mathsf{OGD}}(\boldsymbol{x})\right]\lesssim\sum_{t=1}^{T}\frac{G^{2}}{\mu t}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}t^{\mathsf{p}-1}}\lesssim\frac{G^{2}\left(1+\log T\right)}{\mu}+\frac{\sigma^{\mathsf{p}}D^{2-\mathsf{p}}}{\mu^{\mathsf{p}-1}}\times\begin{cases}T^{2-\mathsf{p}}&\mathsf{p}\in\left(1,2\right)\\ 1+\log T&\mathsf{p}=2\end{cases}.$
@@ Appendix D Missing Proofs for Applications: Nonsmooth Convex Optimization
In this section, we prove the last-iterate convergence for $\mathsf{SGD}$ (i.e., $\mathsf{OGD}$ for stochastic optimization) under heavy-tailed noise. Following the existing literature [22, 25], we will move beyond the standard $G$-Lipschitz convex function considered before and instead focus on a broader class of nonsmooth functions, i.e., $\exists\mu\geq 0$, $G\geq 0$, and $H\geq 0$, such that $G+H>0$ and
$\frac{\mu}{2}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{2}\leq\ell_{t}(\boldsymbol{x})-\ell_{t}(\boldsymbol{y})-\left\langle\nabla\ell_{t}(\boldsymbol{y}),\boldsymbol{x}-\boldsymbol{y}\right\rangle\leq 2G\left\|\boldsymbol{x}-\boldsymbol{y}\right\|+\frac{H}{2}\left\|\boldsymbol{x}-\boldsymbol{y}\right\|^{2},\forall\boldsymbol{x},\boldsymbol{y}\in\mathcal{X},\nabla\ell_{t}(\boldsymbol{y})\in\partial\ell_{t}(\boldsymbol{y}).$ (26)
Notably, such a condition encompasses various commonly studied objectives simultaneously (i.e., strong convexity and smoothness).
Theorem 10 first shows the general last-iterate convergence result, the proof of which is inspired by [25, 50].
@@
Under Assumption 1 (without the third point) and additionally assuming (26) for $\ell_{t}(\boldsymbol{x})=F(\boldsymbol{x})$, for any stepsize $0<\eta_{t}\leq\frac{1}{2H\lor\mu}$ in $\mathsf{OGD}$ (Algorithm 1), let $\gamma_{t}\triangleq\frac{\eta_{t}}{\prod_{s=2}^{t}(1-\mu\eta_{s})}$, we have
$\mathbb{\mathbb{E}}\left[F(\boldsymbol{x}_{T})-F(\boldsymbol{x})\right]\lesssim\frac{(1-\mu\eta_{1})D^{2}}{\sum_{t=1}^{T-1}\gamma_{t}}+G^{2}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}}{\sum_{s=t}^{T-1}\gamma_{s}}+\sigma^{\mathsf{p}}D^{2-\mathsf{p}}\sum_{t=1}^{T-1}\frac{\gamma_{t}\eta_{t}^{\mathsf{p}-1}}{\sum_{s=t}^{T-1}\gamma_{s}},\forall\boldsymbol{x}\in\mathcal{X}.$
@@
$w_{t}\triangleq\frac{\gamma_{T}}{\sum_{s=t}^{T}\gamma_{s}},\forall t\in\left[T\right]$ and $w_{0}\triangleq w_{1}.$ (28)
@@
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\leq\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}}{2\eta_{t}}+\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle-\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}},$
and then bound
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle=\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle+\left\langle\boldsymbol{\epsilon}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$
$\leq$ $F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{2\eta_{t}},$
where, in the second step, we use (5) to bound $\left\langle\boldsymbol{\epsilon}_{t},\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$ (up to change $\eta_{t}$ to $2\eta_{t}$) and the following step to bound $\left\langle\nabla F(\boldsymbol{x}_{t}),\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\rangle$,
[…]
$\leq F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{8\eta_{t}}+\frac{H}{2}\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}$
$\overset{\eta_{t}\leq\frac{1}{2H}}{\leq}F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+8\eta_{t}G^{2}+\frac{3\left\|\boldsymbol{x}_{t}-\boldsymbol{x}_{t+1}\right\|^{2}}{8\eta_{t}}.$
Hence, we know
$\left\langle\boldsymbol{g}_{t},\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\rangle\leq F(\boldsymbol{x}_{t})-F(\boldsymbol{x}_{t+1})+\frac{\left\|\boldsymbol{x}_{t}-\boldsymbol{y}_{t}\right\|^{2}-\left\|\boldsymbol{x}_{t+1}-\boldsymbol{y}_{t}\right\|^{2}}{2\eta_{t}}+8\eta_{t}G^{2}+2^{\mathsf{p}-1}\mathsf{C}(\mathsf{p})\eta_{t}^{\mathsf{p}-1}\left\|\boldsymbol{\epsilon}_{t}\right\|^{\mathsf{p}}D^{2-\mathsf{p}}.$ (30)